# Memory Assisted LLM for Personalized Recommendation System

### v1

@@ 6. Conclusion
In the context of the rapid development of large language models (LLMs), increasing attention has been drawn to the field of personalized LLMs. In this paper, we introduced MAP (Memory-Assisted Personalized LLM), a model that integrates users’ past behaviors while updating with new user requests, demonstrating its advantages in supporting long-term interactions between users and recommender systems in both single-category and cross-domain recommendation tasks. By adopting a retrieval framework to filter and prioritize the most relevant historical data, MAP effectively captures patterns in user preferences, understanding their likes and dislikes over time. To further evaluate MAP in a personalized setting with progressively increasing user history, we designed a new task building upon the LaMP benchmark (Salemi et al., 2023). The proposed MAP structure taps into new potentials for LLM personalization and provides a framework that could guide the design of future personalized LLMs. Our findings underscore the importance of memory-assisted architectures in enhancing both the accuracy and efficiency of personalized recommendations.

### v2

@@ 6. Conclusion
In the context of the rapid development of large language models (LLMs), increasing attention has been drawn to the field of personalized LLMs. In this paper, we introduced MAP (Memory-Assisted Personalized LLM), a model that integrates users’ past behaviors while updating with new user requests, demonstrating its advantages in supporting long-term interactions between users and recommender systems in both single-category and cross-domain recommendation tasks. By adopting a retrieval framework to filter and prioritize the most relevant historical data, MAP effectively captures patterns in user preferences, understanding their likes and dislikes over time. To further evaluate MAP in a personalized setting with progressively increasing user history, we designed a new task based on the LaMP benchmark (Salemi et al., 2023), enabling evaluation based on varying user history sizes and contexts. The proposed MAP structure and the new evaluation task tap into new potentials for personalization of LLM and provide a framework that could guide future work. Our findings underscore the importance of memory-assisted architectures in improving both the accuracy and efficiency of personalized recommendations.